OpenStudy (anonymous):

can someone help me solve 5/6x - 1/6 is less than or equal to x-4

OpenStudy (anonymous):

\[\large \frac{5}{6}x-\frac{1}{6}\leq x-4\]?

OpenStudy (anonymous):

is that a (5/6)x or 5/6x?

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

if you don;t like fractions, multiply both sides by 8 and start with \[\large 5x-1\leq 6x-24\] then it should be much easier

OpenStudy (anonymous):

ooops i means multiply both sides by 6

OpenStudy (anonymous):

thanks